OpenStudy (anonymous):

I need to find an equation of the line containing the given pair of points. (-5,-4) and (-3,-1) Thanks!

OpenStudy (anonymous):

The first thing you need to do is figure out the slope of the line. To do this, use the equation: \[m = (y _{2}-y _{1})\div(x _{2}-x _{1})\] Then, substitute either of your points, and your previously calculated slope, into the equation

OpenStudy (anonymous):

Okay, what if I dont have a slope

OpenStudy (anonymous):

Whoops, sorry... Into the equation \[y-y _{0}=m(x-x _{0})\] and solve for y. \[(x _{0},y _{0})\] You have to calculate the slope. Use the equation I gave you before to solve for m (slope). Use your point (-5,-4) as (x1,y1) and (-3,-1) as (x2,y2).

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

(-5,-4) and (-3,-1) M=(Y1-Y0)/(X1-X0) M= [-1-(-4)]/[-3-(-5)] M= (-1+4)/(-3+5) M=3/2 THEN WE KNOW THAT Y=MX+B USING ONE OF THE POINTS (-3,-1) -1=3/2 (-3) + B -1 = -9/2 +B -1+9/2 = B 7/2 = B THEREFORE THE EQUATION WILL BE: Y= MX+B Y=(3/2)X + 7/2 =)